may discussing d the 90th percentile. Pts, be necessatle a Find core was 64, what is your percentile? 105%: list of number of discharges each day in September, compute the e, and range. Round the mean and median to one decimal point From, median, m mode University Hospital Number of Discharge Days June 20Xx No. of Discharges No. of No. of Day Discharges Day Discharges Day r so that one 63 21 21 42 40 43 50 62 56 62 53 elow which 12 13 62 70 24 25 26 27 28 29 30 27 ution 15 16 17 43 40 28 43 average 63 67 56 19 n a 10
Answer #1

(first part) here number of observation n=30 and sum=1462

therefore mean=sum/n=1462/30=48.73

(second part)median is middle observation, when arrange in ascending order. here there are two middle observation, 15th=50 and 16th=51 so median will be average of these and it is (50+51)/2=50.5

(third part)The mode is a statistical term that refers to the most frequently occurring number found in a set of numbers.

here x=62, repeated maximum 3 times so mode=62

(fouth part)range=maximum-minimum=70-21=49